A mixture of 50.0 g of nitrogen and 35.0 g of oxygen has a total pressure of 0.900 atm. What is the partial pressure of each gas?


1
Expert's answer
2022-05-05T03:39:03-0400

Pi = χi × Ptotal

Pi  - the partial pressure of component i

χi - its mole fraction in the mixture

Ptotal - the total pressure of the mixture


n(N2) = m(N2) / Mr(N2) = 50.0 g / 28 g/mol = 1.79 moles

n(O2) = m(O2) / Mr(O2) = 35 g / 32 g/mol = 1.09 moles

ntotal = n(N2) + n(O2) = 1.79 moles + 1.09 moles = 2.88 moles


χi = number of moles of i / total number of moles


P(N2) = (1.79 mol / 2.88 mol) * 0.900 atm = 0.559 atm

P(O2) = (1.09 mol / 2.88 mol) * 0.900 atm = 0.341 atm